OpenStudy (anonymous):

Describe and correct the error in evaluating the function g(x)=-5x+3 when x=-3 g(-3)=-5(-3)+3 -3g=18 g=-6

OpenStudy (anonymous):

g is not a variable . so this is incorrect: -3g=18

OpenStudy (anonymous):

g(-3) means yo plug x in g(x)=-5x+3 ( x is the independant varaible) ===> g(-3) = -5(-3) +3=15+3=18

OpenStudy (anonymous):

g(-3)=18

OpenStudy (anonymous):

thanks guys appreciate it

hero (hero):

Wow, lol. That's a good way to confuse someone just trying to understand this stuff.

hero (hero):

I can see this as a test question.

OpenStudy (anonymous):

I know lol